Suppose for a given data set the regression equation is: y = 54.78 + 1.45x, and the point (0.00, 20.78) is in the data set. The residual for this point is: A) -25 B) -30 C) -34 D) -64
Answer #1

We have given ,

y = 54.78 + 1.45x

Substitute x=0.00

=>  y = 54.78 + 1.45(0.00)

=54.78..................................................................predicted value for x=0.00

Residual = y-\hat{y}

= 20.78 -54.78

=-34

C) -34